# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ 6.3 Radiation spectrum test
Figure 6: Radiation spectrum test. The top and bottom panels show the distributions of radiation energy density and radiation temperature, respectively, in the steady state. These panels present an xyxy-slice at $z=0$ from the 3D simulation. Dashed lines indicate the surfaces of the clumps, within which the gas temperature is fixed.
7 Astrophysical Application: 3D Simulation of Protostar Evolution in the Early Universe
In this section, we demonstrate an astrophysical application of the radiation hydrodynamics code developed in this study. Specifically, we simulate the evolution of a protostar forming in the early universe under extremely high accretion rates, corresponding to the so-called direct collapse scenario (V. Bromm & A. Loeb, 2003; M. C. Begelman et al., 2006; K. Inayoshi et al., 2014). While a similar problem was previously investigated in K. Kimura et al. (2023), where an artificial Bonnor–Ebert sphere was adopted as the initial condition, here we extend the study by starting from cosmological initial conditions, thereby providing a more self-consistent treatment of the early formation environment.
@@ 7.1 Setup
The approximation of representing the spectrum by a single-temperature Planck distribution introduced in Section 4.3 is reasonable once a protostar has formed and its radiation dominates the system. However, during the earlier stage when the entire cloud is still optically thin and collapsing, such an approximation is not sufficiently accurate because the radiation field has not reached equilibrium and deviates from the Planck distribution. Especially in the collapse phase of a primordial gas cloud, a Planck spectrum at a single effective temperature tends to overestimate the number of ionizing photons with energies above $13.6$ eV. Since the chemical and thermal evolution of primordial gas depends sensitively on the ionization degree, this overproduction of ionizing photons can strongly affect the thermal history. To avoid this, we divide the radiation into two frequency bins separated at $13.6$ eV. For each bin, we evolve both the radiation energy density and photon number density and estimate the spectrum from the mean photon energy, as described in Section 4.3. In computing the mean photon energy with Equation (62) and the opacities with Equations (9), (10), (C7) and (C8), we perform the relevant integrals over the frequency range corresponding to each bin. With this treatment, the radiation naturally approaches a Planck-like spectrum after the protostar forms, while in the optically thin collapsing phase the resulting thermal and chemical evolution agrees with the expected physical behavior (K. Omukai, 2001).
@@
As the initial condition, we adopt the results of a cosmological simulation performed by S. Chon et al. (2016) with the $N$-body and smoothed particle hydrodynamics (SPH) code GADGET2 (V. Springel, 2005). This setup corresponds to the spherical cloud model of S. Chon et al. (2018), who predict that rapid gas accretion at a rate of $\sim 1M_{\odot}~\mathrm{yr}^{-1}$ leads to the formation of a supermassive star exceeding $10^{4}~M_{\odot}$. This formation pathway for supermassive stars is often referred to as the direct collapse scenario, in which the remnant massive black holes provide a promising explanation for the presence of supermassive black holes observed in the early universe. At the initial stage of our simulation, the gas has already become gravitationally unstable due to self-gravity, with a maximum number density of $10^{4}~{\rm cm}^{-3}$. To generate the initial conditions for our grid-based code SFUMATO, we remap SPH data from GADGET2 onto Cartesian grids. In the simulations of S. Chon et al. (2018), high-density regions were replaced with sink particles, and the internal structure of the protostar was not resolved. Instead, protostellar evolution was incorporated as a subgrid model based on 1D stellar evolution calculations by T. Hosokawa et al. (2012, 2013). In contrast, our simulations do not employ sink particles and resolve the protostellar interior directly with much higher spatial resolution. The computational domain spans $\sim 100$ pc. The base grid consists of 64 cells in each direction, and adaptive mesh refinement is applied to ensure that the Jeans length is always resolved with at least 8 cells. The finest spatial resolution achieved in our simulation is $4.7\times 10^{-3}$ au.
@@
The extracted primordial gas cloud is irradiated by strong ultraviolet (UV) radiation from nearby star-forming galaxies, which suppresses H2 formation and triggers gravitational collapse via hydrogen atomic cooling (S. Chon et al., 2018). In our simulations, instead of explicitly including these galaxies within the computational domain, we mimic their effect by imposing a uniform background radiation field with an intensity of $1000~J_{21}$, where the conventional unit $J_{21}$ is defined as $10^{-21}~\mathrm{erg}~\mathrm{s}^{-1}~\mathrm{Hz}^{-1}~\mathrm{cm}^{-2}~\mathrm{sr}^{-1}$. This radiation field induces a high photodissociation rate of H2, thereby preventing molecular hydrogen formation (e.g. K. Sugimura et al., 2014).
7.2 Results
@@
In Figure 7, we present face-on and edge-on views of the protostar at the final snapshot of the simulation, showing the distributions of gas density, temperature, and luminosity. The definition of luminosity is the same as that described in Section 6.2. As seen in the figure, the gas can be divided into two distinct regions. The outer region is characterized by low density and temperature, is optically thin, and exhibits high luminosity. In contrast, the inner region, with a radius of $\sim 4$ au, is dense and hot, optically thick, and shows low luminosity. As shown in the density panels, the gas in the outer region is infalling approximately in free fall with some angular momentum, while in the inner region it exhibits nearly pure rotational motion. The mass of this dense, optically thick region is $\sim 15~M_{\odot}$. From the right panels, the radiation emitted from this region has a luminosity of $\sim 3\times 10^{5}~L_{\odot}$, which is consistent with that of a protostar of the same mass obtained in 1D stellar evolution calculations (T. Hosokawa et al., 2012). In the general context of star formation, the protostar and accretion disk are often regarded as distinct components. In our simulation, however, only a single rotating, optically thick object forms, and no clear boundary exists that separates the protostar from the surrounding accretion disk.
Here we discuss similarities and differences with previous studies. First, our findings are broadly consistent with those of K. Kimura et al. (2023), who started their simulations from an artificial Bonnor–Ebert sphere and likewise found the formation of a single, rotating, optically thick structure. In contrast, Y. Luo et al. (2018), who adopted cosmological initial conditions, reported that the central protostar can be disrupted by gas and radiation pressure, accompanied by strong outflows, after which a new core forms at a different location. Such features are not observed in our present calculation, and the origin of this discrepancy remains to be clarified in future work. We also note that A. Ahmad et al. (2024), in the context of present-day star formation, reported a smooth transition from the protostar to the inner disk, consistent with the structure found in our simulation.
@@
8 Summary and Discussion
@@
The primary motivation for developing the new scheme is to overcome three major limitations inherent to the explicit M1 closure method with the RSLA. First, conventional RSLA treatments inevitably violate total energy conservation in optically thick, radiation-dominated regions, hindering accurate modeling of stellar evolution (Section 3.1). Second, the standard M1 scheme suffers from excessive numerical diffusion in highly optically thick stellar interiors or fails to allow radiation to escape properly across steep optical depth gradients, such as at protostellar surfaces (Section 3.2). Third, most existing RHD simulations employ overly simplified spectral assumptions, which prevent them from following the correct temporal and spatial evolution of the radiation spectrum and lead to inaccurate estimates of protostellar radiative feedback (Section 3.3).
@@
In our new scheme, we overcome the above three limitations as follows. To address the first issue, we introduce a hybrid radiative transfer approach where both RSLA and non-RSLA radiation components are evolved simultaneously (Section 4.1). The non-RSLA component dominates in optically thick regions and restores total energy conservation. For the second issue, we improve the treatment of radiative transfer by redistributing the radiation energy among the streaming, trapped, and non-RSLA components using both local and neighboring optical depth information (Section 4.2). This method allows radiation to diffuse correctly within stellar interiors and to escape properly across steep optical-depth gradients at protostellar surfaces. To resolve the third issue, we evolve photon number densities in addition to the radiation energy densities, which enables us to reconstruct the local spectrum in each cell consistently with the evolving protostar (Section 4.3). This, in turn, allows for a more accurate evaluation of radiative feedback.
@@
We have implemented this new scheme in SFUMATO, a self-gravitating magnetohydrodynamics code with adaptive mesh refinement, which has been widely applied to studies of star formation. The implementation is based on an operator-splitting procedure, in which hydrodynamics, self-gravity, radiation transport, emission and absorption, and momentum exchange are solved step by step in a stable manner (Section 5).
@@
We have validated the method through a series of numerical tests (Section 6), including energy conservation, radiative transfer in optically thick regimes, radiative diffusion across steep optical depth gradients, and reconstruction of evolving radiation spectra. These tests demonstrate that the scheme performs robustly and accurately under conditions relevant to protostellar formation and evolution.
@@
As an application of the new scheme, we have investigated cosmological protostellar evolution in the early universe under the so-called direct collapse scenario, where gas accretion rates are extremely high, within a full cosmological context (Section 7). Our simulations resolve the protostellar interior together with the large-scale anisotropic accretion flow, and demonstrate that the protostar grows while maintaining a smooth connection to its surrounding disk. The overall evolutionary path we obtained is in close agreement with that reported in our previous study K. Kimura et al. (2023), which started from an artificial Bonnor–Ebert sphere. These features cannot be obtained in conventional one-dimensional evolution models, and this application demonstrates that our method can capture protostellar evolution in realistic environments.
A key limitation of our scheme is that the RSLA can artificially reverse the ordering between the gas acoustic signal speed, $v_{\mathrm{sig}}=|\bm{v}|+c_{\mathrm{eff}}$ (where $c_{\mathrm{eff}}$ is the effective sound speed including radiation pressure), and the effective radiation propagation speed $c/\tau$ in the regime $f_{c}c/\tau<v_{\mathrm{sig}}<c/\tau$. However, in the context of protostellar evolution, the optical depth rises sharply across protostellar surfaces and disk surfaces, so the regions where this artificial reversal can occur are confined to narrow layers near their surfaces. Therefore, while this effect is inherent to RSLA-based approaches, we expect it to have only a minor impact on the overall protostellar evolution captured in our simulations.
@@
The versatility of our scheme allows it to be applied to a wide range of astrophysical problems. In the context of protostellar accretion, magnetic fields may play an important role in the region extending from the inner accretion disk to the stellar surface (e.g. S. Takasao et al., 2025; A. C. Mayer et al., 2025). To explore such effects, our framework can be extended toward full radiation magnetohydrodynamics (RMHD) simulations that incorporate magnetic fields self-consistently. The same framework is also relevant to other problems in stellar physics (Y.-F. Jiang, 2023, and references therein), such as mass loss from massive stars, including Wolf–Rayet winds (N. Moens et al., 2022), where the flow direction is opposite to that in protostellar accretion. On even larger spatial scales, our method can be applied to galactic or circumnuclear environments where radiation pressure and gravity interact to produce complex gas dynamics involving both inflows and outflows (T. A. Thompson & T. M. Heckman, 2024, and references therein). Such systems, in which radiation feedback regulates accretion and can even trigger star formation in dense regions, provide another natural target for our scheme. Together, these examples demonstrate the broad applicability of our method and its potential to advance our understanding of radiation-hydrodynamic processes across diverse astrophysical environments.
@@
The authors would like to thank Tomoaki Matsumoto, Kengo Tomida, and Chong-Chong He for fruitful discussions and valuable comments. We also thank Sunmyon Chon for providing the simulation data used in this work and for helpful discussions. TH wishes to express his cordial gratitude to Prof. Takahiro Tanaka for his continuous interest and encouragement. HF wishes to express his heartfelt gratitude to his most revered mentor, Prof. Masayuki Umemura, the Pater Patriae of the Theoretical Astrophysics Group at the University of Tsukuba, as well as to Prof. Ken Ohsuga, for their constant interest, valuable advice, and encouragement. The numerical simulations were carried out on the XC50 (Aterui II) and XC2000 (Aterui III) at the Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan, through the courtesy of Prof. Eiichiro Kokubo, and on the Yukawa-21 at the Yukawa Institute for Theoretical Physics, Kyoto University. This research was supported in part by Grants-in-Aid for Scientific Research ($K$: 24KJ0015, KS: 22KK0043, 24H00002, TH: 19KK0353, 22H00149, HF: 23K13139, KO: 22H00149) from the Japan Society for the Promotion of Science, and by the Astrobiology Center Project research (HF: AB0511). TH also appreciates the financial support from ISHIZUE 2024 provided by Kyoto University and the Kyoto University Foundation.

### v2

@@ VI.3 Radiation spectrum test
Figure 6: Radiation spectrum test. The top and bottom panels show the distributions of radiation energy density and radiation temperature, respectively, in the steady state. These panels present an xyxy-slice at $z=0$ from the 3D simulation. Dashed lines indicate the surfaces of the clumps, within which the gas temperature is fixed. The simulation is performed in a cubic box of size $320$ au with a uniform grid of $256^{3}$.
VI.4 Computational overhead of RT
We comment on the increase in wall-time of RT introduced by the newly implemented scheme. We measure the wall-time fractions for the test calculation in Section VI.3, as summarized in Table 5. The computational cost increases due to the inclusion of the non-RSLA component. However, because this diffusive RT involves a simpler flux evaluation than the M1 RT, its associated overhead is limited to $8\%$, compared to $41\%$ for the M1 RT. The evaluation of $\tau_{\mathrm{cell}}$, which requires collecting information from neighboring cells, is a relatively simple operation and accounts for only $1\%$ of the total wall-time. Evolving the photon number density accounts for $50\%$ of the total RT wall-time, effectively doubling the computational cost of the RT solver, because the procedures for evolving the photon number density are similar to those for evolving the photon energy density. Overall, the total computational cost of the newly implemented scheme is approximately a factor of $100/41\simeq 2.4$ higher than that of a M1 RT scheme evolving only the radiation energy density.
Table 5: Wall-time percentages for each computational operation in the radiation spectrum test in Section VI.3.
Operation percentage [%]
M1 RT of RSLA component 41
Diffusive RT of Non-RSLA component 8
Estimation of $\tau_{\mathrm{cell}}$ 1
RT of photon number density 50
VII Astrophysical Application: 3D Simulation of Protostar Evolution in the Early Universe
In this section, we demonstrate an astrophysical application of the radiation hydrodynamics code developed in this study. Specifically, we simulate the evolution of a protostar forming in the early universe under extremely high accretion rates, corresponding to the so-called direct collapse scenario (Bromm and Loeb, 2003; Begelman et al., 2006; Inayoshi et al., 2014). While a similar problem was previously investigated in Kimura et al. (2023), where an artificial Bonnor–Ebert sphere was adopted as the initial condition, here we extend the study by starting from cosmological initial conditions, thereby providing a more self-consistent treatment of the early formation environment.
@@ VII.1 Setup
The approximation of representing the spectrum by a single-temperature Planck distribution introduced in Section IV.3 is reasonable once a protostar has formed and its radiation dominates the system. However, during the earlier stage when the entire cloud is still optically thin and collapsing, such an approximation is not sufficiently accurate because the radiation field has not reached equilibrium and deviates from the Planck distribution. Especially in the collapse phase of a primordial gas cloud, a Planck spectrum at a single effective temperature tends to overestimate the number of ionizing photons with energies above $13.6$ eV. Since the chemical and thermal evolution of primordial gas depends sensitively on the ionization degree, this overproduction of ionizing photons can strongly affect the thermal history. To avoid this, we divide the radiation into two frequency bins separated at $13.6$ eV. For each bin, we evolve both the radiation energy density and photon number density and estimate the spectrum from the mean photon energy, as described in Section IV.3. In computing the mean photon energy with Equation (62) and the opacities with Equations (9), (10), (C7) and (C8), we perform the relevant integrals over the frequency range corresponding to each bin. With this treatment, the radiation naturally approaches a Planck-like spectrum after the protostar forms, while in the optically thin collapsing phase the resulting thermal and chemical evolution agrees with the expected physical behavior (Omukai, 2001).
@@
As the initial condition, we adopt the results of a cosmological simulation performed by Chon et al. (2016) with the $N$-body and smoothed particle hydrodynamics (SPH) code GADGET2 (Springel, 2005). This setup corresponds to the spherical cloud model of Chon et al. (2018), who predict that rapid gas accretion at a rate of $\sim 1M_{\odot}~\mathrm{yr}^{-1}$ leads to the formation of a supermassive star exceeding $10^{4}~M_{\odot}$. This formation pathway for supermassive stars is often referred to as the direct collapse scenario, in which the remnant massive black holes provide a promising explanation for the presence of supermassive black holes observed in the early universe. At the initial stage of our simulation, the gas has already become gravitationally unstable due to self-gravity, with a maximum number density of $10^{4}~{\rm cm}^{-3}$. To generate the initial conditions for our grid-based code SFUMATO, we remap SPH data from GADGET2 onto Cartesian grids. In the simulations of Chon et al. (2018), high-density regions were replaced with sink particles, and the internal structure of the protostar was not resolved. Instead, protostellar evolution was incorporated as a subgrid model based on 1D stellar evolution calculations by Hosokawa et al. (2012, 2013). In contrast, our simulations do not employ sink particles and resolve the protostellar interior directly with much higher spatial resolution. The computational domain spans $\sim 100$ pc. The base grid consists of 64 cells in each direction, and adaptive mesh refinement is applied to ensure that the Jeans length is always resolved with at least 8 cells. The finest spatial resolution achieved in our simulation is $4.7\times 10^{-3}$ au.
@@
The extracted primordial gas cloud is irradiated by strong ultraviolet (UV) radiation from nearby star-forming galaxies, which suppresses H2 formation and triggers gravitational collapse via hydrogen atomic cooling (Chon et al., 2018). In our simulations, instead of explicitly including these galaxies within the computational domain, we mimic their effect by imposing a uniform background radiation field with an intensity of $1000~J_{21}$, where the conventional unit $J_{21}$ is defined as $10^{-21}~\mathrm{erg}~\mathrm{s}^{-1}~\mathrm{Hz}^{-1}~\mathrm{cm}^{-2}~\mathrm{sr}^{-1}$. This radiation field induces a high photodissociation rate of H2, thereby preventing molecular hydrogen formation (e.g. Sugimura et al., 2014).
VII.2 Results
@@
In Figure 7, we present face-on and edge-on views of the protostar at the final snapshot of the simulation, showing the distributions of gas density, temperature, and luminosity. The definition of luminosity is the same as that described in Section VI.2. As seen in the figure, the gas can be divided into two distinct regions. The outer region is characterized by low density and temperature, is optically thin, and exhibits high luminosity. In contrast, the inner region, with a radius of $\sim 4$ au, is dense and hot, optically thick, and shows low luminosity. As shown in the density panels, the gas in the outer region is infalling approximately in free fall with some angular momentum, while in the inner region it exhibits nearly pure rotational motion. The mass of this dense, optically thick region is $\sim 15~M_{\odot}$. From the right panels, the radiation emitted from this region has a luminosity of $\sim 3\times 10^{5}~L_{\odot}$, which is consistent with that of a protostar of the same mass obtained in 1D stellar evolution calculations (Hosokawa et al., 2012). In the general context of star formation, the protostar and accretion disk are often regarded as distinct components. In our simulation, however, only a single rotating, optically thick object forms, and no clear boundary exists that separates the protostar from the surrounding accretion disk.
Here we discuss similarities and differences with previous studies. First, our findings are broadly consistent with those of Kimura et al. (2023), who started their simulations from an artificial Bonnor–Ebert sphere and likewise found the formation of a single, rotating, optically thick structure. In contrast, Luo et al. (2018), who adopted cosmological initial conditions, reported that the central protostar can be disrupted by gas and radiation pressure, accompanied by strong outflows, after which a new core forms at a different location. Such features are not observed in our present calculation, and the origin of this discrepancy remains to be clarified in future work. We also note that Ahmad et al. (2024), in the context of present-day star formation, reported a smooth transition from the protostar to the inner disk, consistent with the structure found in our simulation.
@@
VIII Discussion
VIII.1 Local break down of energy conservation in optically thick system
Although we introduce the Non-RSLA component to achieve energy conservation in the optically thick system in Section IV.1, energy conservation locally breaks down in optically thick systems if they contain regions where $\tau_{\mathrm{cell}}<2/(3f_{c})$, since the RSLA component $E^{\mathrm{R}}_{\mathrm{rad}}$ dominates in such regions. This issue can become significant in extremely high-resolution simulations, where $\tau_{\mathrm{cell}}$ may fall below $2/(3f_{c})$ even within stellar interiors. As discussed in Section III.1, when energy is transferred from the gas to the radiation field, only a fraction $f_{c}$ of the transferred energy is actually received by the radiation. In an extreme case where the gas energy is entirely converted into radiation, the RSLA reduces the total energy to only $f_{c}$ times its correct value. At this time, because the radiation energy density scales as $T^{4}$, this reduction corresponds to a temperature deviation of $f_{c}^{1/4}$.
However, in typical RHD star formation simulations, where the Jeans length is resolved by several tens of cells, $\tau_{\mathrm{cell}}$ remains sufficiently large throughout most of the stellar interior, and the regions with $\tau_{\mathrm{cell}}<2/(3f_{c})$ are confined to narrow layers near the stellar surface. As a result, the impact of the above effect on the global energy budget is expected to be small. To illustrate this, consider representative photospheric conditions of the Sun: $T\simeq 5.8\times 10^{3}\ {\rm K}$, $\rho\simeq 10^{-9}\ {\rm g\,cm^{-3}}$, and a Rosseland–mean opacity $\kappa_{\mathrm{0R}}\simeq 1.5\times 10^{-2}\ {\rm cm^{2}\,g^{-1}}$. The corresponding Jeans length is $\lambda_{\mathrm{J}}\approx c_{\mathrm{s}}\sqrt{\pi/(G\rho)}\simeq 1.8\times 10^{14}\ {\rm cm}$ while its optical depth is enormous, $\tau_{\mathrm{J}}\sim\kappa_{\mathrm{0R}}\,\rho\,\lambda_{\mathrm{J}}\simeq 2.7\times 10^{3}$. Thus, even if the Jeans scale were resolved by several tens of cells, each cell would still remain optically thick ($\tau_{\rm cell}\gg 1$) even near the photosphere. Furthermore, as the density increases rapidly toward the stellar interior, $\tau_{\rm cell}$ also rises steeply and exceeds $2/(3f_{c})$ near the surface.
VIII.2 Radiation behavior in the transition region between the RSLA and non-RSLA components
Figure 8: Transition test between the Non-RSLA and RSLA components. The blue solid and orange dashed lines represent the spherically averaged radial profile of physical radiation energy density and conserved radiation energy density in the steady-state from our simulation. The green dotted line denotes the analytic solution. The gray dashed and dotted lines indicate the radius at which $\tau_{\mathrm{cell}}=2/(3f_{c})$ and the computational cell size ($3.13$ au). The simulation is performed in a cubic box of size $200$ au with a uniform grid of $64^{3}$.
In our new scheme, the radiation field is decomposed into the RSLA and non-RSLA components according to Equation (49) as mentioned in Section IV.1. With this decomposition, we can solve radiative transfer using the RSLA component in the optically thin limit and the non-RSLA component in the optically thick limit. However, in the intermediate regime, the radiation field is constructed by interpolating between the optically thin and optically thick limits. In this section, we examine how the radiation field behaves in this transition region. We perform the calculation in which the radiation is injected at the center of the computational box and transported outward, using the same setup as in Section VI.1.2, except for the spatial resolution and the gas density distribution. For the resolution, the computational box width is $200$ au, and we resolve the box with $64^{3}$ cells, corresponding to a cell size of $3.13~\mathrm{au}$. For the gas density distribution, we adopt an exponential gas density profile,
$\rho(r)=A\exp(-r/B)~\mathrm{g~cm^{-3}},$ (103)
where $r$ is the radial distance measured in centimeters, $A=2.7\times 10^{18}$ and $B=9\times 10^{13}$ cm. As a result, the optical depth per cell $\tau_{\mathrm{cell}}$ is $10^{7}$ at the center of the computational box and decreases to unity at the outer boundary. We set $f_{c}=10^{3}$, so that the transition between the dominance of the non-RSLA and RSLA components occurs at $\tau_{\mathrm{cell}}=2/(3f_{c})\simeq 667$, which corresponds to a radius of $60~\mathrm{au}$. In this case, assuming a steady state, we can obtain an analytic solution in the same manner as in Section VI.1.2,
$E_{\mathrm{rad}}=\frac{3\sigma_{0}^{\nu}AL}{4\pi c}\left[\frac{e^{-r/B}}{r}+\frac{1}{B}\,\mathrm{Ei}\!\left(-\frac{r}{B}\right)\right],$ (104)
where $\mathrm{Ei}$ denotes the exponential integral. In the calculation, we impose Dirichlet boundary conditions using the value given by Equation (104).
Figure 8 shows the spherically averaged radial profiles of the physical radiation energy density, $f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$ (the blue solid line), and the conserved radiation energy density, $E^{\mathrm{con}}_{\mathrm{rad}}=E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$ (the orange dashed line), in the steady state. For comparison, the analytic solution is shown (the green dotted line). We also indicate the radius at which $\tau_{\mathrm{cell}}=2/(3f_{c})$ and the cell size by the gray dashed and dotted lines, respectively. Near the center ($r<3.13~\mathrm{au}$), a small discrepancy between the simulation result and the analytic solution is present, which arises from the finite size of the injection cell, as discussed in Section VI.1.2.
In Figure 8, we show the conserved radiation energy density $E^{\mathrm{con}}_{\mathrm{rad}}$ by the orange dashed line. In the inner region ($r<30~\mathrm{au}$), where the non-RSLA component is dominant and $E^{\mathrm{con}}_{\mathrm{rad}}\simeq E^{\mathrm{N}}_{\mathrm{rad}}$, $E^{\mathrm{con}}_{\mathrm{rad}}$ follows the analytic solution. In contrast, in the outer region ($r>80~\mathrm{au}$), where the RSLA component dominates and $E^{\mathrm{con}}_{\mathrm{rad}}\simeq E^{\mathrm{R}}_{\mathrm{rad}}$, $E^{\mathrm{con}}_{\mathrm{rad}}$ follows an analytic solution elevated by a factor of $1/f_{c}$. This enhancement arises because, when the RSLA is effective, the radiation propagation speed is suppressed, leading to a larger accumulation of radiation energy in a steady state. The orange dashed line in Figure 8 connects these inner and outer solutions.
The physical radiation energy density, $f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$ (the blue solid line in Figure 8), agrees well with the analytical solution in both the inner and outer regions. In the outer region where RSLA-component $E^{\mathrm{R}}_{\mathrm{rad}}$ is dominant, the physical radiation energy density is approximated as $f_{c}E^{\mathrm{con}}_{\mathrm{rad}}$ and consistent with the analytic solution. Then, the physical radiation energy density gradually approaches $E^{\mathrm{con}}_{\mathrm{rad}}$ toward smaller radii since the fraction of non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$ becomes larger. In the intermediate region, however, $E^{\mathrm{con}}_{\mathrm{rad}}$ has not yet fully converged to the analytic solution, which results in a deviation of the physical radiation energy density from the analytic solution. At sufficiently small radii, $E^{\mathrm{con}}_{\mathrm{rad}}$ converges to the analytic solution, and consequently the physical radiation energy density also agrees with it.
As demonstrated in this test, our scheme successfully reproduces the physically correct state in both the inner and outer regions. However, a discrepancy from the physically correct state remains in the transition region between the non-RSLA and RSLA components. For the present application focusing on protostellar evolution, this transition layer is confined to an extremely thin region near the protostellar surface, and this discrepancy is therefore not expected to have a significant impact on our results. In fact, in Section VI.2, we demonstrate that our scheme can accurately follow radiative transport across regions where the optical depth changes abruptly. Moreover, this discrepancy can be mitigated by choosing the reduction factor $f_{c}$ closer to unity.
VIII.3 Choice of reduction factor $f_{c}$
In our scheme, $f_{c}$ is a free parameter. The choice of $f_{c}$ is typically guided by the requirement that the ordering between the gas acoustic signal speed $v_{\mathrm{sig}}=|\bm{v}|+c_{\mathrm{eff}}$ (where $c_{\mathrm{eff}}$ is the effective sound speed including radiation pressure) and the effective radiation propagation speed $\tilde{c}/\max(\tau,1)$ should not be reversed relative to the ordering in the non-RSLA system (Skinner and Ostriker, 2013). For the protostellar evolution problems considered in this work, we choose $f_{c}$ such that the correct ordering is maintained in the optically thin regions outside the protostar, while allowing a reversal of the ordering in the dense interior of the protostar. Generally, when a simulation is designed to resolve both the exterior and interior of a protostar, it is not possible to satisfy this ordering criterion throughout the entire computational domain. This is because the effective radiation propagation speed is larger in the optically thin exterior, whereas the gas effective sound speed is larger in the dense interior. As a result, for any choice of $f_{c}$, the RSLA inevitably introduces an artificial reversal of these characteristic speeds somewhere between the exterior and interior. However, because the optical depth rises extremely steeply across the protostellar surface, our choice ensures that any such artificial reversal is confined to narrow layers near the surface, and we therefore expect it to have only a minor influence on the global protostellar evolution captured in our simulations.
IX Summary
@@
The primary motivation for developing the new scheme is to overcome three major limitations inherent to the explicit M1 closure method with the RSLA. First, conventional RSLA treatments inevitably violate total energy conservation in optically thick, radiation-dominated regions, hindering accurate modeling of stellar evolution (Section III.1). Second, the standard M1 scheme suffers from excessive numerical diffusion in highly optically thick stellar interiors or fails to allow radiation to escape properly across steep optical depth gradients, such as at protostellar surfaces (Section III.2). Third, most existing RHD simulations employ overly simplified spectral assumptions, which prevent them from following the correct temporal and spatial evolution of the radiation spectrum and lead to inaccurate estimates of protostellar radiative feedback (Section III.3).
@@
In our new scheme, we overcome the above three limitations as follows. To address the first issue, we introduce a hybrid radiative transfer approach where both RSLA and non-RSLA radiation components are evolved simultaneously (Section IV.1). The non-RSLA component dominates in optically thick regions and restores total energy conservation. For the second issue, we improve the treatment of radiative transfer by redistributing the radiation energy among the streaming, trapped, and non-RSLA components using both local and neighboring optical depth information (Section IV.2). This method allows radiation to diffuse correctly within stellar interiors and to escape properly across steep optical-depth gradients at protostellar surfaces. To resolve the third issue, we evolve photon number densities in addition to the radiation energy densities, which enables us to reconstruct the local spectrum in each cell consistently with the evolving protostar (Section IV.3). This, in turn, allows for a more accurate evaluation of radiative feedback.
@@
We have implemented this new scheme in SFUMATO, a self-gravitating magnetohydrodynamics code with adaptive mesh refinement, which has been widely applied to studies of star formation. The implementation is based on an operator-splitting procedure, in which hydrodynamics, self-gravity, radiation transport, emission and absorption, and momentum exchange are solved step by step in a stable manner (Section V).
@@
We have validated the method through a series of numerical tests (Section VI), including energy conservation, radiative transfer in optically thick regimes, radiative diffusion across steep optical depth gradients, and reconstruction of evolving radiation spectra. These tests demonstrate that the scheme performs robustly and accurately under conditions relevant to protostellar formation and evolution.
@@
As an application of the new scheme, we have investigated cosmological protostellar evolution in the early universe under the so-called direct collapse scenario, where gas accretion rates are extremely high, within a full cosmological context (Section VII). Our simulations resolve the protostellar interior together with the large-scale anisotropic accretion flow, and demonstrate that the protostar grows while maintaining a smooth connection to its surrounding disk. Such a continuous star–disk interface implies that entropy is not generated abruptly at the stellar surface but is supplied gradually through dissipative processes within the disk, which may affect its radial inflation. The overall evolutionary path we obtained is in close agreement with that reported in our previous study Kimura et al. (2023), which started from an artificial Bonnor–Ebert sphere. These features cannot be obtained in conventional one-dimensional evolution models, and this application demonstrates that our method can capture protostellar evolution in realistic environments.
@@
The versatility of our scheme allows it to be applied to a wide range of astrophysical problems. In the context of protostellar accretion, magnetic fields may play an important role in the region extending from the inner accretion disk to the stellar surface (e.g. Takasao et al., 2025; Mayer et al., 2025). To explore such effects, our framework can be extended toward full radiation magnetohydrodynamics (RMHD) simulations that incorporate magnetic fields self-consistently. The same framework is also relevant to other problems in stellar physics (Jiang, 2023, and references therein), such as mass loss from massive stars, including Wolf–Rayet winds (Moens et al., 2022), where the flow direction is opposite to that in protostellar accretion. On even larger spatial scales, our method can be applied to galactic or circumnuclear environments where radiation pressure and gravity interact to produce complex gas dynamics involving both inflows and outflows (Thompson and Heckman, 2024, and references therein). Such systems, in which radiation feedback regulates accretion and can even trigger star formation in dense regions, provide another natural target for our scheme. Together, these examples demonstrate the broad applicability of our method and its potential to advance our understanding of radiation-hydrodynamic processes across diverse astrophysical environments.
@@
The authors would like to thank Tomoaki Matsumoto, Kengo Tomida, and Chong-Chong He for fruitful discussions and valuable comments. We also thank Sunmyon Chon for providing the simulation data used in this work and for helpful discussions. TH wishes to express his cordial gratitude to Prof. Takahiro Tanaka for his continuous interest and encouragement. HF wishes to express his heartfelt gratitude to his most revered mentor, Prof. Masayuki Umemura, the Pater Patriae of the Theoretical Astrophysics Group at the University of Tsukuba, as well as to Prof. Ken Ohsuga, for their constant interest, valuable advice, and encouragement. The numerical simulations were carried out on the XC50 (Aterui II) and XD2000 (Aterui III) at the Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan, through the courtesy of Prof. Eiichiro Kokubo, and on the Yukawa-21 at the Yukawa Institute for Theoretical Physics, Kyoto University. This research was supported in part by Grants-in-Aid for Scientific Research ($K$: 24KJ0015, KS: 22KK0043, 24H00002, TH: 19KK0353, 22H00149, HF: 23K13139, KO: 22H00149) from the Japan Society for the Promotion of Science, and by the Astrobiology Center Project research (HF: AB0511). TH also appreciates the financial support from ISHIZUE 2024 provided by Kyoto University and the Kyoto University Foundation.
Data Availability
The input files used for the numerical tests and simulation data are available from the corresponding author upon reasonable request.